OpenStudy (jazzyluvsu123):

Miranda is making cookies. The recipe makes 18 cookies and uses 2/3 of sugar. SHe needs to make 3 dozen cookies. How much sugar will Miranda need ? I need to set it up as a proportion @satellite73 @mathmale or @mathstudent55 PLZ HELP FREE MEDAL !!!

OpenStudy (anonymous):

3 dozen is \(3\times 12=36\) which is twice \(18\) so she needs twice as much sugar

OpenStudy (anonymous):

double it in other words twice \(\frac{2}{3}\) is \(\frac{4}{3}\)

OpenStudy (jazzyluvsu123):

so 1&1/3 cups of sugar ? @satellite73

OpenStudy (jazzyluvsu123):

@satellite73 plz help mee

OpenStudy (anonymous):

yes

OpenStudy (jazzyluvsu123):

I got 4/3 but that is not on my answer sheet

OpenStudy (jazzyluvsu123):

yay I got it right

OpenStudy (anonymous):

\[1\tfrac{1}{3}\] is right if that is a choice

OpenStudy (jazzyluvsu123):

thanks I got 6 more questions to go !!

OpenStudy (anonymous):

jeez it is endless

OpenStudy (jazzyluvsu123):

i know right

OpenStudy (anonymous):

be done soon enough i am sure